# Decision Quotient: A Regime-Sensitive Complexity Theory of Exact Relevance Certification

### 16 Mar 2026 v1

@@ 9.4.1 Witness-Checking Lower Bound
Theorem 9.18 (Witness-Checking Lower Bound).
@@
Let $S=\{0,1\}^{n}$. Empty-set sufficiency is exactly the claim that Opt is constant on all states.
@@
9.4.2 Approximation Gap on the Shifted Family
@@
One coordinate acts as a gate: toggling it changes the optimizer even before the formula variables are considered, so optimum size is never 0. If $\varphi$ is a tautology, every branch behind that gate behaves identically, so all non-gate coordinates become irrelevant and the singleton gate coordinate is sufficient. If $\varphi$ is not a tautology, choose a falsifying assignment $a$. For each formula coordinate $i$, compare the open reference state with the state that inserts $a$ at coordinate $i$: these states agree on all other coordinates but induce different optimal-action sets, so coordinate $i$ is relevant. Together with gate relevance, this forces every coordinate to be present in every sufficient set. That is what creates the exact $1$ versus $n+1$ gap.
Theorem 9.19 (Approximation-hardness theorem on the shifted family).
@@
Tautology $\Rightarrow$ threshold passes. If $\varphi$ is a tautology, the optimum support size on the shifted family is $1$, realized by the singleton gate coordinate. A factor-$\rho$ solver therefore returns a sufficient set of size at most $\rho\cdot 1=\rho$, so $|\mathcal{A}(\varphi)|\leq\rho$.
@@
Theorem 9.20 (Counted-runtime threshold decider).
@@
9.4.3 Set-Cover Transfer Boundary
Theorem 9.21 (Exact set-cover equivalence on the gadget family).
@@ 9.4.3 Set-Cover Transfer Boundary
If $I$ is not a cover, choose an uncovered universe element $u$. Then every coordinate in $I$ takes the same value on $(\mathsf{false},u)$ and $(\mathsf{true},u)$, but their optimal-action sets differ. So $I$ is not sufficient.
If $I$ is a cover, then for every universe element there exists some coordinate in $I$ that separates the two tagged states attached to that element. Hence two states that agree on all coordinates in $I$ cannot disagree in the tag component in a way that changes the optimizer. The optimizer therefore factors through the $I$-projection, so $I$ is sufficient.
@@
Corollary 9.22 (Conditional approximation transfer).
@@
9.5 Externalized Relevance and Simplicity Tax
The final implication concerns architectural compression. The main point is interpretive rather than algebraic: shrinking a central interface does not remove exact relevance; it relocates the unresolved part of that relevance to some other part of the system. The simplicity tax is the name for that externalized burden.
Definition 9.23 (Relevant Support).
For decision problem $\mathcal{D}$, let
[…]
as in Definition 2.8.
Definition 9.24 (Intrinsic and Central Degrees of Freedom).
For a model $M$ with native coordinate set $A_{M}\subseteq\{1,\ldots,n\}$:
$\mathrm{intrinsicDOF}(\mathcal{D}):=|R(\mathcal{D})|,\qquad\mathrm{centralDOF}(M,\mathcal{D}):=|R(\mathcal{D})\cap A_{M}|.$
Definition 9.25 (Simplicity Tax).
$\mathrm{simplicityTax}(M,\mathcal{D}):=|R(\mathcal{D})\setminus A_{M}|.$
The following partition identity makes this precise: any exact simplification that removes relevant coordinates from the central interface must discharge their effect somewhere else in the system.
Proposition 9.26 (Simplicity-Tax Identity).
For every finite-coordinate pair $(M,\mathcal{D})$,
$\mathrm{centralDOF}(M,\mathcal{D})+\mathrm{simplicityTax}(M,\mathcal{D})=\mathrm{intrinsicDOF}(\mathcal{D}).$
(L: HD5, DQ9.)
Proof.
Immediate from the partition $R(\mathcal{D})=(R(\mathcal{D})\cap A_{M})\;\dot{\cup}\;(R(\mathcal{D})\setminus A_{M})$.  $\blacksquare$
In other words, central simplification changes where exact relevance is handled, not how much exact relevance the deployment must ultimately account for.
Proposition 9.27 (Linear Externalization Under a Per-Site Cost Model).
If each unresolved relevant coordinate induces one unit of per-site external work, then for $N$ decision sites:
$\mathrm{ExternalWork}(N)=N\cdot\mathrm{simplicityTax}(M,\mathcal{D}).$
(L: HD26.)
Proof.
Each site pays one unit for each unresolved relevant coordinate, so the total cost is the number of sites times the number of unresolved relevant coordinates.  $\blacksquare$
Theorem 9.28 (Amortization Threshold).
Let $H_{\mathrm{central}}>0$ be one-time centralization cost and $\lambda>0$ the per-site cost coefficient. If $\mathrm{simplicityTax}(M,\mathcal{D})>0$, then for
$N>\frac{H_{\mathrm{central}}}{\lambda\cdot\mathrm{simplicityTax}(M,\mathcal{D})},$
the complete model is cheaper than repeated external handling:
$H_{\mathrm{central}}<\lambda N\cdot\mathrm{simplicityTax}(M,\mathcal{D}).$
(L: HD4.)
Proof.
Repeated external handling costs $\lambda N\cdot\mathrm{simplicityTax}(M,\mathcal{D})$, while the complete model costs $H_{\mathrm{central}}$ once. Solving the strict inequality $H_{\mathrm{central}}<\lambda N\cdot\mathrm{simplicityTax}(M,\mathcal{D})$ for $N$ gives the stated threshold.  $\blacksquare$
Corollary 9.29 (Hardness Conservation).
Fix a model $M$ and decision problem $\mathcal{D}$. Coordinates in $R(\mathcal{D})\setminus A_{M}$ are still decision-relevant for exact behavior preservation. Hence any exact deployment must handle them somewhere: by enlarging the central model, by performing local resolution, by demanding extra queries, or by abstaining on some cases. In the hard exact regime, Theorem 9.18 and Theorem 9.14 imply that this burden cannot be discharged for free by a polynomial-budget exact certifier.
(L: DQ9, HD23.)
@@ 9.5 Externalized Relevance and Simplicity Tax
So omitted relevant coordinates must be handled somewhere outside the central interface. In the hard exact regime, the witness lower bound shows that exact certification can require exponentially many witness checks, and the integrity-resource theorem shows that polynomial-budget exact competence is unavailable under $\textsf{P}\neq\textsf{coNP}$. Therefore this external burden cannot, in general, be eliminated for free by a polynomial-budget exact certifier.  $\blacksquare$
10 Related Work
10.1 Formalized Complexity Theory and Mechanized Reductions
@@
10.2 Rough Sets, Reducts, and Attribute Reduction
@@
10.3 Backdoor Tractability in CSPs
A closely related line of work studies backdoor tractability for constraint satisfaction problems [33, 2]. In that framework, a backdoor is a minimal variable set whose fixation collapses a CSP to a tractable class (e.g., one admitting a majority polymorphism). Bessiere et al. [2] showed that testing membership in tractable classes characterized by majority or conservative Malt’sev polymorphisms is polynomial-time, but finding a minimal backdoor to such a class is W[2]-hard when the tractable subset is unknown and fixed-parameter tractable only when that subset is given as input. This directly parallels our structural results: coordinate sufficiency in decision problems and backdoor tractability in CSPs both formalize the same intuition—that discovering what matters is fundamentally harder than exploiting a known structure. Our bounded-treewidth reduction (Theorem 8.13) explicitly connects to their CSP algorithms and backdoor methodology.
@@ 10.3 Backdoor Tractability in CSPs
The key structural difference is this: general decision tables can admit multiple incomparable minimal reducts, necessitating a combinatorial search through the reduct lattice. Under the exact optimal-action preservation predicate studied here, Proposition 2.9 proves there is exactly one minimal sufficient set: the relevant-coordinate set itself. This uniqueness is the structural reason for the complexity collapse. The search layer vanishes, reducing exact minimization to parallel coNP relevance checks. To our knowledge, this specific structural collapse, from $\Sigma_{2}^{P}$ search to coNP relevance containment driven by optimizer-quotient uniqueness, is a novel classification not captured by general rough-set bounds.
@@
Proposition 10.1 (Static Sufficiency as Reduct Preservation for the Induced Decision Table).
Given a finite decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$, form the induced decision table whose condition attributes are the coordinates $X_{i}$ and whose decision label at state $s$ is the set-valued class label $\text{Opt}(s)\subseteq A$. Then a coordinate set $I$ is sufficient in the sense of this paper if and only if the projection to $I$ preserves the decision label of the induced table. Consequently, the minimal sufficient sets of $\mathcal{D}$ are exactly the reducts of this induced decision table.
@@
By definition, $I$ is sufficient exactly when any two states agreeing on $I$ have the same optimal-action set. In the induced decision table, the decision label of state $s$ is precisely $\text{Opt}(s)$, so the same condition says that projection to $I$ preserves the decision label. Minimality on the present side is therefore exactly reduct minimality for the induced table.  $\blacksquare$
Rough-set reducts are formulated relative to indiscernibility relations or decision tables, whereas the object here is the optimizer map of a decision problem and the induced exact preservation of optimal-action sets. Proposition 10.1 shows that the static regime can be recast as reduct preservation for the induced decision table. What the present paper adds is: (i) the collapse theorem showing minimum sufficiency collapses to relevance containment, bypassing the general NP-hardness barrier for exact minimization; and (ii) the systematic extension to stochastic and sequential regimes, which takes the reduct question beyond the classical rough-set setting. The stochastic and sequential layers introduce conditional-comparison and temporal-structure phenomena that are outside the usual rough-set formulation.
@@
10.4 Decision-Theoretic Sufficiency, Informativeness, and Value of Information
There is also a strong conceptual connection to statistical decision theory and the theory of informativeness. Classical sufficient statistics, value-of-information analysis, and comparison of experiments ask when one information structure is at least as useful for decision making as another [3, 29, 26, 15]. That tradition studies when information can be compressed, compared, or discarded without harming decisions. The optimizer quotient and the exact sufficiency predicate in this paper should be read against that background.
At the same time, the present work studies a different layer of the problem. The focus here is not statistical estimation, asymptotic identifiability, or expected utility under a fixed experiment class. It is the computational certification problem of proving that a coordinate projection preserves the optimal-action correspondence exactly. The connection to classical decision theory is therefore one of ancestry and motivation rather than direct theorem transfer. The paper owes that literature recognition for posing, in older language, the same high-level question: what information is genuinely needed for a decision?
10.5 Abstraction, Bisimulation, and Homomorphisms in Planning and Reinforcement Learning
@@
That literature is structurally close to the present paper and should be credited as such. The main difference is that our predicate is coordinate sufficiency for preserving optimal-action sets under an explicit coordinate-hiding operation. We do not study arbitrary abstract state maps, approximate bisimulation metrics, or value-function approximation guarantees. Instead, we ask for exact certification that a chosen coordinate interface is enough. This exact-certification emphasis is what the paper adds beyond prior abstraction results: the question is not only whether a smaller representation can support good control, but whether one can certify that no omitted coordinate changes the optimal-action correspondence. The contribution here is therefore not abstraction theory in general, but a unified exact-certification account in which static, stochastic, and sequential regimes are analyzed as variants of one decision-relevance question.
10.6 Explanations, Anchors, and Minimal Decision Rules
The paper should also acknowledge adjacent work in explainable AI and local rule-based explanations, especially anchor-style explanations [28]. Readers will notice the word “anchor” and map it to that literature. The resemblance is real: both settings isolate a restricted condition under which a local piece of information suffices to support a decision or prediction. More broadly, explanation methods based on local rules, counterfactuals, or minimal sufficient subsets share the same informal ambition of identifying a small decision-preserving core.
The difference is again exactness and scope. Anchor explanations in XAI are typically local, model-agnostic, approximate, or precision-calibrated. Our anchor-sufficiency problem is an exact certification problem inside a fully specified decision model, and its complexity is governed by an existential choice followed by a universal fiberwise verification condition. So the paper is not an explanation method, but it is in direct conversation with explanation work that asks what small local structure is enough to preserve a decision.
There is an even closer logical neighbor in recent work on reasons behind decisions, prime-implicant style explanations, and exact decision justifications [7]. This shows that exact small decision-preserving structures are a live contemporary topic, not only a classical one. That literature is concerned with exact structural reasons for a decision outcome and therefore lies much nearer to the present paper than generic post-hoc explanation methods do. The overlap is substantial at the conceptual level: both ask for exact small structures that preserve a decision. The difference is that our objects are coordinate sufficiency, minimum sufficiency, and anchor sufficiency inside coordinate-structured decision problems, and the paper’s main contribution is a regime-sensitive complexity classification of these predicates rather than a logical semantics of reasons alone.
10.7 Abduction, Diagnosis, and Exact Explanatory Cores
The minimum and anchor queries share structural kinship with abduction and diagnosis frameworks, which study explanatory hypotheses sufficient for a target consequence. The anchor query retains an outer existential choice analogous to explanatory discovery, while the inner verification condition preserves optimizer structure. This logical-complexity intuition motivates the $\Sigma_{2}^{P}$-completeness of static anchor sufficiency, which remains hard even when minimum sufficiency collapses to coNP.
10.8 Causality, Responsibility, and Structural Explanation
The paper also belongs in conversation with structural accounts of causality, explanation, and responsibility [25, 31, 13, 6]. A relevant coordinate is precisely a coordinate whose perturbation can change the decision, and the optimizer quotient isolates exactly the distinctions that matter for optimal-action structure. The paper focuses on exact decision preservation under coordinate projection and on the computational cost of certifying that preservation, whereas causality and responsibility frameworks typically emphasize semantic characterization or intervention-level explanation.
10.9 Oracle, Query, and Access-Based Lower Bounds

### 31 Mar 2026 v2

@@ X-C1 Witness-Checking Lower Bound
Theorem X.5footnotemark: footnotetext: : WD1-3 (Witness-Checking Lower Bound).
@@
Let $S=\{0,1\}^{n}$. Empty-set sufficiency is exactly the claim that $\mathrm{Opt}$ is constant on all states.
@@
X-C2 Approximation Gap on the Shifted Family
@@
One coordinate acts as a gate: toggling it changes the optimizer even before the formula variables are considered, so optimum size is never 0. If $\varphi$ is a tautology, every branch behind that gate behaves identically, so all non-gate coordinates become irrelevant and the singleton gate coordinate is sufficient. If $\varphi$ is not a tautology, choose a falsifying assignment $a$. For each formula coordinate $i$, compare the open reference state with the state that inserts $a$ at coordinate $i$: these states agree on all other coordinates but induce different optimal-action sets, so coordinate $i$ is relevant. Together with gate relevance, this forces every coordinate to be present in every sufficient set. That is what creates the exact 11 versus $n+1$ gap.
Theorem X.6footnotemark: footnotetext: : DQ20, DQ39, DQ43 (Approximation-hardness theorem on the shifted family).
@@
Tautology $\Rightarrow$ threshold passes. If $\varphi$ is a tautology, the optimum support size on the shifted family is 11, realized by the singleton gate coordinate. A factor-$\rho$ solver therefore returns a sufficient set of size at most $\rho\cdot 1=\rho$, so $|\mathcal{A}(\varphi)|\leq\rho$.
@@
Theorem X.7footnotemark: footnotetext: : DQ23-24 (Counted-runtime threshold decider).
@@
X-C3 Set-Cover Transfer Boundary
Theorem X.8footnotemark: footnotetext: : DQ11, DQ34 (Exact set-cover equivalence on the gadget family).
@@ X-C3 Set-Cover Transfer Boundary
If II is not a cover, choose an uncovered universe element $u$. Then every coordinate in II takes the same value on $(\mathsf{false},u)$ and $(\mathsf{true},u)$, but their optimal-action sets differ. So II is not sufficient.
If II is a cover, then for every universe element there exists some coordinate in II that separates the two tagged states attached to that element. Hence two states that agree on all coordinates in II cannot disagree in the tag component in a way that changes the optimizer. The optimizer therefore factors through the II-projection, so II is sufficient.
@@
Corollary X.9footnotemark: footnotetext: : DQ25, DQ32-33 (Conditional approximation transfer).
@@
X-D Externalized Relevance and Simplicity Tax
The final implication concerns architectural compression. Compression relocates relevance, it does not remove it. When coordinates relevant for exact decisions are omitted from a central model, the unresolved burden must be handled elsewhere in the system: by local resolution, extra queries, or abstention. The simplicity tax is the name for that externalized burden.
Corollary X.10footnotemark: footnotetext: : DQ9, HD23 (Hardness Conservation).
Fix a model $M$ and decision problem $\mathcal{D}$. Coordinates in $R(\mathcal{D})\setminus A_{M}$ are still decision-relevant for exact behavior preservation. Hence any exact deployment must handle them somewhere: by enlarging the central model, by performing local resolution, by demanding extra queries, or by abstaining on some cases. In the hard exact regime, Theorem IX.6 implies that this burden cannot be discharged for free by a polynomial-budget exact certifier.
@@ X-D Externalized Relevance and Simplicity Tax
So omitted relevant coordinates must be handled somewhere outside the central interface. In the hard exact regime, Theorem IX.6 shows that polynomial-budget exact competence is unavailable under $\mathrm{P}\neq\mathrm{coNP}$. Therefore this external burden cannot, in general, be eliminated for free by a polynomial-budget exact certifier.  $\blacksquare$
The simplicity tax operationalizes the certification trilemma (Corollary 7.3) for architectural compression. When exact relevance cannot be certified, systems cannot simply discard coordinates without consequence: the unresolved burden is conserved across the interface and must be paid elsewhere. This principle connects the complexity map to practical deployment constraints: the cost of exact certification is either paid at the center through computational investment, or distributed to the sites as externalized handling.
X-E Artifact Scope and the Finite-to-Asymptotic Boundary
The Lean 4 artifact verifies the finite combinatorial core used throughout Sections 3–5: reduction correctness with size bounds, explicit-state deciders and step-counted searches, bridge lemmas, and witness/checking schemas used by stochastic upper-bound arguments. The artifact checks the finite mechanisms—reduction constructions, witness structures, explicit-state deciders—that underpin the paper’s asymptotic complexity claims. The asymptotic class placements (coNP-completeness, $P$-hardness, PSPACE-completeness, NPˆPP membership) are argued in the text using standard complexity-theoretic conventions, as is typical for the field.
The paper-level layer is the asymptotic lift: oracle-machine formulations, class memberships, and ETH-transfer arguments. In particular, the artifact does not formalize oracle Turing machines or polynomial-space machine semantics; those are argued in the manuscript using standard complexity-theoretic conventions.
[…]
Summary of what is verified vs. what is assumed.
Claim Verified in Lean Paper-level
Reduction correctness Polynomial-size output, truth-value preservation
Explicit-state decidability Exact algorithm with step bound
Witness/checking pattern Finite existential/universal structure Oracle-machine encoding
coNP-completeness TAUTOLOGY $\to$ empty-set sufficiency (finite) Asymptotic class membership
$\mathsf{PP}$-hardness MAJSAT $\to$ stochastic decisiveness (finite) Majority-vote characterization
$\mathsf{PSPACE}$-completeness TQBF $\to$ sequential sufficiency (finite) Polynomial-space TM theory
$\mathsf{NP}^{\mathsf{PP}}$ upper bound Witness/checking schema Oracle machine construction
ETH lower bound Reduction to empty-set sufficiency Asymptotic reduction chain
TABLE VI: Summary of what is verified in Lean vs. what is argued at paper-level.
The finite core is solid; the asymptotic lifts are standard but not mechanized.
XI Related Work
XI-A Formalized Complexity Theory and Mechanized Reductions
@@
XI-B Rough Sets, Reducts, and Attribute Reduction
@@
XI-C Backdoor Tractability in CSPs
A closely related line of work studies backdoor tractability for constraint satisfaction problems [14, 2]. In that framework, a backdoor is a minimal variable set whose fixation collapses a CSP to a tractable class (e.g., one admitting a majority polymorphism). Bessiere et al. [2] showed that testing membership in tractable classes characterized by majority or conservative Malt’sev polymorphisms is polynomial-time, but finding a minimal backdoor to such a class is W[2]-hard when the tractable subset is unknown and fixed-parameter tractable only when that subset is given as input. This directly parallels our structural results: coordinate sufficiency in decision problems and backdoor tractability in CSPs both formalize the same intuition, that discovering what matters is fundamentally harder than exploiting a known structure. Our bounded-treewidth reduction (Theorem VIII.13) explicitly connects to their CSP algorithms and backdoor methodology. Subsequent work in parameterized complexity has rigorously bounded the tractability of discovering these minimal structural sets across varying graph widths and backdoor types [15].
@@ XI-C Backdoor Tractability in CSPs
The key structural difference is this: general decision tables can admit multiple incomparable minimal reducts, necessitating a combinatorial search through the reduct lattice. Under the exact optimal-action preservation predicate studied here, Proposition II.9 proves there is exactly one minimal sufficient set: the relevant-coordinate set $R(\mathcal{D})$ itself. This uniqueness follows from the optimizer quotient structure: states in the same quotient class share the same optimal-action set, so any sufficient set must separate all quotient classes, and the relevant-coordinate set is precisely the minimal set achieving this separation.
The complexity consequence is significant. In general rough-set reduct computation, multiple minimal reducts can exist, placing minimal-reduct search in $\Sigma_{2}^{P}$. Here, uniqueness collapses the search problem: finding a minimal sufficient set reduces to checking which coordinates are relevant (coNP-complete). The $\Sigma_{2}^{P}$ search layer vanishes because there is no need to search among incomparable candidates; the single minimal set is uniquely determined by relevance. This structural collapse—from $\Sigma_{2}^{P}$ search to coNP relevance containment—is driven by optimizer-quotient uniqueness and is, to our knowledge, a novel classification not captured by general rough-set bounds.
@@
Proposition XI.1footnotemark: footnotetext: : DP1-2 (Static Sufficiency as Reduct Preservation for the Induced Decision Table).
Given a finite decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$, form the induced decision table whose condition attributes are the coordinates $X_{i}$ and whose decision label at state $s$ is the set-valued class label $\mathrm{Opt}(s)\subseteq A$. Then a coordinate set II is sufficient in the sense of this paper if and only if the projection to II preserves the decision label of the induced table. Consequently, the minimal sufficient sets of $\mathcal{D}$ are exactly the reducts of this induced decision table.
@@
By definition, II is sufficient exactly when any two states agreeing on II have the same optimal-action set. In the induced decision table, the decision label of state $s$ is precisely $\mathrm{Opt}(s)$, so the same condition says that projection to II preserves the decision label. Minimality on the present side is therefore exactly reduct minimality for the induced table.  $\blacksquare$
Rough-set reducts are formulated relative to indiscernibility relations or decision tables, whereas the object here is the optimizer map of a decision problem and the induced exact preservation of optimal-action sets. Proposition XI.1 shows that the static regime can be recast as reduct preservation for the induced decision table. What the present paper adds is: (i) the collapse theorem showing minimum sufficiency collapses to relevance containment, bypassing the general NP-hardness barrier for exact minimization; and (ii) the systematic extension to stochastic and sequential regimes, which takes the reduct question beyond the classical rough-set setting. The stochastic and sequential layers introduce conditional-comparison and temporal-structure phenomena that are outside the usual rough-set formulation.
@@
XI-D Decision-Theoretic Sufficiency, Informativeness, and Value of Information
Classical decision theory and informativeness ask when one information structure is at least as useful as another [20, 21, 22, 23]. Our setting is the computational certification layer of that question: exact coordinate projection preserving optimal-action correspondence.
XI-E Abstraction, Bisimulation, and Homomorphisms in Planning and Reinforcement Learning
@@
That literature is structurally close to the present paper and should be credited as such. The main difference is that our predicate is coordinate sufficiency for preserving optimal-action sets under an explicit coordinate-hiding operation. We do not study arbitrary abstract state maps, approximate value-function guarantees, or the quantitative bisimulation metrics pioneered by Ferns et al. [29] to measure behavioral similarity. While recent work has successfully leveraged such metrics to learn robust, task-invariant representations in high-dimensional deep reinforcement learning (e.g., Zhang et al. [30]), our focus remains strictly on the combinatorial complexity of exact certification. Instead of optimizing a continuous representation space to discard task-irrelevant distractors, we ask whether one can exactly certify that no omitted coordinate changes the optimal-action correspondence. The contribution here is therefore not abstraction theory in general, but a unified exact-certification account in which static, stochastic, and sequential regimes are analyzed as variants of one decision-relevance question.
XI-F Explanatory and Structural Reasoning Links
Anchor-style explanations, abductive explanations, and exact reason/justification frameworks all study small structures sufficient for a decision outcome [31, 32, 33, 34]. Our contribution is not a new explanation semantics but a regime-sensitive complexity classification for exact coordinate-preservation predicates.
The same applies to causality/responsibility traditions [35, 36, 37, 38]: they motivate relevance notions, while this paper studies the algorithmic certification cost of exact preservation under coordinate hiding.
XI-G Oracle, Query, and Access-Based Lower Bounds